Several layers of glass are stacked together each layer reduces the light passing through it by 5%
A) what percent of the light passes through 10 layers of glass?
Thanks :)

Answer by ikleyn(52921)

.
Several layers of glass are stacked together each layer reduces the light passing through it by 5%
A) what percent of the light passes through 10 layers of glass?
~~~~~~~~~~~~~~~~

0.95   = 95% passes throuhg 1 (one) layer.


0.95^2 = 0.9025 = 90.25% passes through 2 (two) layears.


0.95^3 = 0.8574 = 85.74% passes through 3 (three) layears.

. . . . . . . . . . . .


0.95^10 = 0.5984 = 59.84% passes through 10 (ten) layears.    ANSWER

Solved.